How much would be in your savings account in eight years after depositing $240 today if the bank pays 8 percent per year? (Do not d cales a Round your rower to 2 decmai piace%)
Answer #1

FV = PV(1+r)^n

FV = 240(1+0.08)^8

FV = 240(1.85093021)

FV = 444.22325

Therefore Future Value = 444.22